# Uncertainty Quantification with the Empirical Neural Tangent Kernel

### v2

@@ H.7 Additional VMSP Results
We use this section to expand the empirical evaluation of NUQLS. In the top left of Figure˜6, we observe that on a ResNet50 model trained on SVHN, with CIFAR-10 as the OoD test data, NUQLS outperforms all other methods. In the top right of Figure˜6, we evaluated NUQLS on the ImageNet dataset in order to display the scalability of our method to larger datasets. We employed the pre-trained weights for a ResNet50, as found on torch.hub, and used the ImageNet-o dataset as our OOD test set (Hendrycks et al., 2021). Due to resource budget constraints, we were only able to compare NUQLS against a baseline method, though future research could include a comparison of other methods against NUQLS, using VMSP, on ImageNet. As can be seen, we see excellent separation between correct and incorrect predictions, while only adequate separation between correct and OOD points. We note from Bitterwolf et al. (2023) that approximately $21\%$ of the images in ImageNet-o are actually in-distribution; hence, we see that NUQLS correctly quantifies the variance of these test sets. We also provide comparison of NUQLS with additional competing methods. In the bottom left of Figure˜6, we compare NUQLS against the BE method. Note that BE requires modification to the structure of a neural network; hence, we compare BE on a modified WideResNet-34-1, of which an implementation of the correct modifications existed (Franchi et al., 2023), trained on FashionMNIST. We again observe that NUQLS shows greater separation than BE for VMSP across correct predictions versus incorrect or OOD predictions. Finally, in the bottom right of Figure˜6, we compare NUQLS against further competing methods, specifically BDE, SGLD (Welling & Teh, 2011) (which we include to demonstrate the inferior performance of MCMC methods in this regime) and SNGP on ResNet9 trained on FMNIST (note SNGP requires modifications to the training procedure). We see that BDE performs similarly to DE, and SLGD performs similarly to SWAG, which is unsurprising considering their respective constructions. We also observe that NUQLS still performs the strongest across all methods using the VMSP violin plots. The respective median and skew values for Figure˜6 are displayed in Table˜10 and Table˜11; these values also evidence the strong performance of NUQLS.

### v3

@@ H.7 Additional VMSP Results
We use this section to expand the empirical evaluation of NUQLS. In the top left of Figure˜6, we observe that on a ResNet50 model trained on SVHN, with CIFAR-10 as the OoD test data, NUQLS outperforms all other methods. In the top right of Figure˜6, we evaluated NUQLS on the ImageNet dataset in order to display the scalability of our method to larger datasets. We employed the pre-trained weights for a ResNet50, as found on torch.hub, and used the ImageNet-o dataset as our OOD test set (Hendrycks et al., 2021). Due to resource budget constraints, we were only able to compare NUQLS against a baseline method, though future research could include a comparison of other methods against NUQLS, using VMSP, on ImageNet. As can be seen, we see excellent separation between correct and incorrect predictions, while only adequate separation between correct and OOD points. We note from Bitterwolf et al. (2023) that approximately $21\%$ of the images in ImageNet-o are actually in-distribution; hence, we see that NUQLS correctly quantifies the variance of these test sets. We also provide comparison of NUQLS with additional competing methods. In the bottom left of Figure˜6, we compare NUQLS against the BE method. Note that BE requires modification to the structure of a neural network; hence, we compare BE on a modified WideResNet-34-1, of which an implementation of the correct modifications existed (Franchi et al., 2023), trained on FashionMNIST (with MNIST as OOD). We again observe that NUQLS shows greater separation than BE for VMSP across correct predictions versus incorrect or OOD predictions. Finally, in the bottom right of Figure˜6, we compare NUQLS against further competing methods, specifically BDE, SGLD (Welling & Teh, 2011) (which we include to demonstrate the inferior performance of MCMC methods in this regime) and SNGP on ResNet9 trained on FMNIST (with MNIST as OOD); note SNGP requires modifications to the training procedure. We see that BDE performs similarly to DE, and SLGD performs similarly to SWAG, which is unsurprising considering their respective constructions. We also observe that NUQLS still performs the strongest across all methods using the VMSP violin plots. The respective median and skew values for Figure˜6 are displayed in Table˜10 and Table˜11; these values also evidence the strong performance of NUQLS.